Subtonic7:

Convert the complex number from rectangular form to polar form. Show all work for full credit. z=-14+8i

Subtonic7:

Ai is dumb i swear Yes, the complex number -14 + 8i can definitely be represented in polar form. To convert the complex number from rectangular form (a + bi) to polar form (r∠θ), where r is the magnitude and θ is the angle: First, we calculate the magnitude (r) using the formula: r = √(a^2 + b^2) In this case: a = -14 b = 8 r = √((-14)^2 + (8)^2) r = √(196 + 64) r = √260 r = 2√65 Next, we calculate the angle (θ) using the formula: θ = tan^(-1)(b/a) In this case: θ = tan^(-1)(8 / -14) θ ≈ -0.5191 radians Therefore, the polar form of the complex number -14 + 8i is: z = 2√65∠-0.5191

ImJayOkay:

z = 2√65∠-0.5191

Luigi0210:

So you are given z, which is depicted as: \[\Large z=x+yi\] And you need to get it into this form: \[\Large z=r(cos\theta+isin\theta)\] In order to do that, you're gonna need to find r, and r is: \[\Large r = |z|= \sqrt{x^2+y^2}\] Plug in the values of x and y

ImJayOkay:

plug in

Subtonic7:

@imjayokay wrote:
plug in
die

ImJayOkay:

@subtonic7 wrote:
@imjayokay wrote:
plug in
die
ok

babymay06:

this is easy, you plug in the numbers to see if the equation is right. 🤦🏾‍♀️

Subtonic7:

@luigi0210 wrote:
So you are given z, which is depicted as: \[\Large z=x+yi\] And you need to get it into this form: \[\Large z=r(cos\theta+isin\theta)\] In order to do that, you're gonna need to find r, and r is: \[\Large r = |z|= \sqrt{x^2+y^2}\] Plug in the values of x and y
welp thanks guys

ImJayOkay:

@babymay06 wrote:
this is easy, you plug in the numbers to see if the equation is right. 🤦🏾‍♀️
Well explained answero

babymay06:

@imjayokay wrote:
@babymay06 wrote:
this is easy, you plug in the numbers to see if the equation is right. 🤦🏾‍♀️
Well explained answero
gracias wae 😌💯

Subtonic7:

@imjayokay wrote:
z = 2√65∠-0.5191
Wrong

ImJayOkay:

@subtonic7 wrote:
@imjayokay wrote:
z = 2√65∠-0.5191
Wrong
It correct

babymay06:

@subtonic7 wrote:
@imjayokay wrote:
z = 2√65∠-0.5191
Wrong
🤦🏾‍♀️

Subtonic7:

@imjayokay wrote:
@subtonic7 wrote:
@imjayokay wrote:
z = 2√65∠-0.5191
Wrong
It correct
THATS WHAT THE AI SAID 💀

Subtonic7:

@subtonic7 wrote:
Ai is dumb i swear Yes, the complex number -14 + 8i can definitely be represented in polar form. To convert the complex number from rectangular form (a + bi) to polar form (r∠θ), where r is the magnitude and θ is the angle: First, we calculate the magnitude (r) using the formula: r = √(a^2 + b^2) In this case: a = -14 b = 8 r = √((-14)^2 + (8)^2) r = √(196 + 64) r = √260 r = 2√65 Next, we calculate the angle (θ) using the formula: θ = tan^(-1)(b/a) In this case: θ = tan^(-1)(8 / -14) θ ≈ -0.5191 radians Therefore, the polar form of the complex number -14 + 8i is: z = 2√65∠-0.5191
@imjayokay wrote:
z = 2√65∠-0.5191

ImJayOkay:

@subtonic7 wrote:
@imjayokay wrote:
@subtonic7 wrote:
@imjayokay wrote:
z = 2√65∠-0.5191
Wrong
It correct
THATS WHAT THE AI SAID 💀
no it not

Subtonic7:

BRUH

ImJayOkay:

righ

Subtonic7:

WRONG

surjithayer:

\[z=-14+8\iota\] compare with \[z=r(cos \theta+\iota sin \theta)\] \[r cos \theta=-14\] \[r sin \theta=8\] square and add \[r^2(cos^2 \theta+sin^2 \theta=(-14)^2+8^2=196+64=260=4 \times 65\] \[r=2 \sqrt{65}\] again divide \[tan \theta=-8/14=-4/7\] \[\theta=\tan ^{-1}(-4/7)=-29.74488 \approx -30 ^\circ\] sin is positive cos is negative so theta lies in second quadrant. so \theta=180-30=150 \[z=2 \sqrt{65}(150^\circ)\]

surjithayer:

\[z=2\sqrt{65}e ^{150 ^\circ~\iota}\]

surjithayer:

\[z=2\sqrt{65}(\cos 150+\iota \sin 150)\]